value of the mass of synthetic gypsum produced for every kilogram of sulcus dioxide recycled, given that the gypsum can be sold to a gyproc wallboard factory for 75 cents per kilogram


1
Expert's answer
2022-01-25T17:04:03-0500

According to the reaction, one kilogram of SO2 equeals:

n (SO2) = m(SO2)/M(SO2) = 1,000 g / 64 g/mol = 15.625 mol,

where n - number of moles, m - mass, M - molecular weight.

As a result, the same the same number of moles of CaSO4 · 2H2O(s) (synthetic gypsum) was produced (n (CaSO4 · 2H2O) = 15.625 mol), its mass equals:

m (CaSO4 · 2H2O) = n (CaSO4 · 2H2O) × M (CaSO4 · 2H2O) = 15.625 mol × 172 g/mol = 2,687.5 g = 2.7 kg.

This amount of gypsum can be sold for:

2.7 x 0.75 = 2.03 dollars